# When isometry and equivalence for skew constacyclic codes coincide

### v1

@@ 5.3. An interesting correction
Let $\sigma\in{\rm Aut}(S)$ have order $\mathbf{m}$, fixed field $S_{0}$, and let $a,b,\alpha\in S^{\times}$. If there is a Hamming-weight preserving isomorphism $G:S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-a)\to S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$ with $G|_{S}=\tau$, where $\tau\in\operatorname{Aut}(S)$ commutes with $\sigma$, then
@@
Suppose there exists a Hamming weight preserving algebra isomorphism given by $G:S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-a)\to S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$ such that $G|_{S}=\tau$ and $G(t)=\alpha t^{k}$ for some $\alpha\in S^{\times}$, $0<k\leq\mathbf{n}-1$.
@@
    Suppose $G_{\tau,\alpha,k}:S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-a)\to S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$ is an isomorphism with $k>1$ for some $\alpha\in S^{\times}$. Then $k\equiv 1\mod\mathbf{m}$, $\gcd(k,\mathbf{n})=1$, and $G_{\tau,\alpha}:S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-a)\to S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b^{k})$ is also an isomorphism.
@@
    Then $G_{\tau,\alpha}:S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-a)\to S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$ is an isomorphism if and only if $G_{id,\alpha}:S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-\tau(a))\to S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$ is an isomorphism.
@@
    Also, for any $a\in\mathrm{Fix}(\tau)=\{b\in S:\tau(b)=b\}$, we have $G_{\operatorname{id},\alpha}:S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-a)\to S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$ is an isomorphism if and only if $G_{\tau,\alpha}:S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-a)\to S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$ is an isomorphism.
@@
$G_{\tau,\alpha,k}:S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-a)\to S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$
@@
$G_{\tau,\alpha}:S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-a)\to S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b^{k})$
@@
$G_{id,\alpha}:S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-\tau(a))\to S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$
being a $S_{0}$-algebra isomorphism.
@@
Moreover, (a) implies that two classes of skew constacyclic codes that are $(\mathbf{n},\sigma)$-isometric and related to principal left ideals in $S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-a)$, respectively to principal left ideals in $S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$, will generate two classes of codes in $S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-a^{k})$, respectively, in $S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$, that are $(m,\sigma)$-equivalent.
@@
We derive the answer for the case that $S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-a)$ is nonassociative, that is, when $(\mathbf{n},\sigma)$-equivalence and $(\mathbf{n},\sigma)$-isometry coincide. Note that we do not limit ourselves to those $a$ such that $t^{n}-a$ is reducible.
@@
Suppose $K=\mathbb{F}_{p^{r}}$ and $\sigma(x)=x^{p^{s}}$ with $s|r$ so that $\mathbf{m}=r/s$ and $K_{0}=\mathbb{F}_{p^{s}}$. When $a\notin K_{0}$, $K[t;\sigma]/K[t;\sigma](t^{\mathbf{n}}-a)$ is nonassociative, and the number of different $(\mathbf{n},\sigma)$-isometry classes is $N$ where
$N=\begin{cases}\gcd([\mathbf{n}]_{s},p^{r}-1)&\text{if $\mathbf{m}\nmid\mathbf{n}$; and}\\ \gcd([\mathbf{n}]_{s},p^{r}-1)-1&\text{if $\mathbf{m}|\mathbf{n}$.}\end{cases}$
Moreover, if $w=\gcd([\mathbf{n}]_{s},p^{r}-1)$ and $t=\gcd([\mathbf{n}]_{s},\frac{p^{r}-1}{p^{s}-1})$ then the number of $b\in aN_{\mathbf{n}}^{\sigma}(\mathbb{F}_{p^{r}}^{\times})$ such that $b\notin\mathbb{F}_{p^{s}}$ is
  • $(p^{r}-1)/w$ if $\mathbf{m}|\mathbf{n}$ and $a\notin\mathbb{F}_{p^{s}}$
  • 0 if $\mathbf{m}|\mathbf{n}$ and $a\in\mathbb{F}_{p^{s}}$
  • $(p^{r}-1)/w$ if $\mathbf{m}\nmid\mathbf{n}$ and $a=\xi^{A}$ such that $t\nmid A$
  • $(p^{r}-1)/w-t$ if $\mathbf{m}\nmid\mathbf{n}$ and $a=\xi^{A}$ and $t|A$.
Note that when $\mathbf{m}|\mathbf{n}$, we have $N_{\mathbf{n}}^{\sigma}(K)\subset K_{0}$, and so if $a\in bN_{\mathbf{n}}^{\sigma}(K)$ and $a\notin K_{0}$ then $b\notin K_{0}$. Thus the second case corresponds to the isometry classes of associative Petit algebras.
@@
The $(\mathbf{n},\sigma)$-isometry class of $(\sigma,a)$-constacyclic codes is represented by all $b\in\mathbb{F}_{p^{r}}\setminus\mathbb{F}_{p^{s}}$ such that
@@
for some $\alpha\in K^{\times}$, that is, if and only if $a$ and $b$ lie in the same coset of the subgroup $N_{\mathbf{n}}^{\sigma}(K^{\times})$. Let $\xi$ be a primitive element of $K$; it has order $p^{r}-1$. Then $N_{\mathbf{n}}^{\sigma}(K^{\times})$ is generated by $N_{\mathbf{n}}^{\sigma}(\xi)=\xi^{[\mathbf{n}]_{s}}=\xi^{w}$ with $w=\gcd([\mathbf{n}]_{s},p^{r}-1)$, and thus the order of this group is
@@
The number of equivalence classes of codes is the number of cosets of this group, yielding
$\gcd([\mathbf{n}]_{s},p^{r}-1).$
However, not all of these will be nonassociative so this is an upper bound.
For a more precise count, choose $a\in\mathbb{F}_{p^{r}}\setminus\mathbb{F}_{p^{s}}$ and write $a=\xi^{A}$ for some integer $A$. Then we have $b\in aN_{\mathbf{n}}^{\sigma}(\mathbb{F}_{p^{r}}^{\times})$ if and only if
$b\in\{\xi^{A+kw}\mid 0\leq k<\frac{p^{r}-1}{w}\}.$
Recall that $\xi^{h}\in\mathbb{F}_{q^{s}}$ if and only if $(p^{r}-1)|h(p^{s}-1)$ or $h\equiv 0\mod\frac{p^{r}-1}{p^{s}-1}$. Thus we need to subtract from this coset all choices of $k$ such that
(5.1) $A+kw\equiv 0\mod\frac{p^{r}-1}{p^{s}-1}.$
Recall $\mathbf{m}=r/s$. If $\mathbf{m}|\mathbf{n}$ or $r|s\mathbf{n}$ then $p^{r}-1$ divides $p^{sn}-1$, so that $(p^{r}-1)/(p^{s}-1)$ divides $\gcd([\mathbf{n}]_{s},p^{r}-1)=w$. Thus (5.1) holds if and only if $A\equiv 0\mod(p^{r}-1)/(p^{s}-1)$, that is, if and only if $a\in\mathbb{F}_{p^{s}}$ already. This gives the first two statements.
Now set $t=\gcd(w,\frac{p^{r}-1}{p^{s}-1})=\gcd([\mathbf{n}]_{s},\frac{p^{r}-1}{p^{s}-1})=\gcd(\frac{p^{s\mathbf{n}}-1}{p^{s}-1},\frac{p^{r}-1}{p^{s}-1}).$ So $t$ divides $w$ and $\frac{p^{r}-1}{p^{s}-1}$ so if $t\nmid A$ then (5.1) has no solution, meaning all elements of the coset are outside of $\mathbb{F}_{p^{s}}$.
On the other hand, if $t|A$ write $A=\ell t$; then (5.1) holds for all $k$ such that $\ell+k(w/t)\equiv 0\mod\frac{p^{r}-1}{t(p^{s}-1)}$. Since $\gcd(w/t,\frac{p^{r}-1}{t(p^{s}-1)})=1,$ this has a unique solution modulo $\frac{p^{r}-1}{t(p^{s}-1)}$, hence $t$ solutions modulo $\frac{p^{r}-1}{(p^{s}-1)}$. Therefore $t$ elements of the coset lie in $\mathbb{F}_{p^{s}}$, yielding the final statement.
Therefore, if $\mathbf{m}|\mathbf{n}$ we have exactly $\gcd([\mathbf{n}]_{s},p^{r}-1)-1$ of these $(\mathbf{n},\sigma)$ isometry (or $\mathbf{n},\sigma)$-equivalence) classes, since every coset except one is fully contained in $\mathbb{F}_{p^{r}}\setminus\mathbb{F}_{p^{s}}$.
On the other hand, if $\mathbf{m}\nmid\mathbf{n}$ then every coset meets $\mathbb{F}_{p^{s}}\setminus\mathbb{F}_{p^{r}}$, so we have $w$ of these $(\mathbf{n},\sigma)$ isometry (or $\mathbf{n},\sigma)$-equivalence) classes. ∎
@@
At one extreme, suppose that $N_{\mathbf{n}}^{\sigma}(S^{\times})=S^{\times}$. Then all $S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-a)$ are $(\mathbf{n},\sigma)$-equivalent, and in particular all are $(\mathbf{n},\sigma)$-equivalent to $S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-1)$, which corresponds to the class of skew cyclic codes. Moreover, equivalence and $(\mathbf{n},\sigma)$-equivalence coincide in this case. When $\mathbf{m}\nmid\mathbf{n}$, then the notions of equivalence, isometry and $(\mathbf{n},\sigma)$-isometry coincide as well. Hence for all $a\in S^{\times}$, the skew $(\sigma,a)$-constacyclic codes of length $\mathbf{n}$ over $S$ are equivalent to the skew cyclic codes of length $\mathbf{n}$.
As a special case, let $K=\mathbb{F}_{p^{r}}$, $\sigma(x)=x^{p^{s}}$ with $s|r$ and $K_{0}=\mathbb{F}_{p^{s}}$. If $\gcd([\mathbf{n}]_{s},p^{r}-1)=1$ then $N_{\mathbf{n}}^{\sigma}(K^{\times})=K^{\times}$ so all $K[t;\sigma]/K[t;\sigma](t^{\mathbf{n}}-a)$ are both $(\mathbf{n},\sigma)$-equivalent, as was already observed in [OaNA25, Corollary 2] and [BBB21, Proposition 1].
@@
At the other extreme, when $\gcd([\mathbf{n}]_{s},p^{r}-1)=p^{r}-1$ (as in Example 5.2), then $N_{\mathbf{n}}^{\sigma}(K^{\times})=\{1\}$ and so no two distinct $a,b\in K^{\times}$ will be $(\mathbf{n},\sigma)$-equivalent. Thus, there are a plethora of distinct skew constacyclic codes. That said, there always at least $\mathbf{m}$ choices for $\tau$ (corresponding to $\sigma^{\ell}$ for each $1\leq\ell<\mathbf{m}$).
@@
On the other hand, when $S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-a)$ is associative we can prove a powerful equivalence (inspired by the flavour of [OaNA25, Theorem 4]) that holds for any commutative ring $S$ and any $\tau$ commuting with $\sigma$.
@@
    $K[t;\sigma]/K[t;\sigma](t^{\mathbf{n}}-a)$ and $K[t;\sigma]/K[t;\sigma](t^{\mathbf{n}}-b)$ are isometric via $G_{\tau,\alpha,k}$;
@@
    $K[t;\sigma]/K[t;\sigma](t^{\mathbf{n}}-b^{k})$ and $K[t;\sigma]/K[t;\sigma](t^{\mathbf{n}}-a)$ are equivalent via $G_{\tau,\alpha}$, where $k$ satisfies $k\equiv 1\mod\mathbf{m}$ and $\gcd(k,\mathbf{n})=1$.
@@
Corollary 5.10.
Suppose $K=\mathbb{F}_{p^{r}}$ and $\sigma(x)=x^{p^{s}}$ so that $\mathbf{m}=r/s$. Suppose that $K[t;\sigma]/K[t;\sigma](t^{\mathbf{n}}-a)$ is associative, that is, $a\in K_{0}^{\times}$ and $\mathbf{m}|\mathbf{n}$. Then the number of different $(\mathbf{n},\sigma)$-equivalence classes of skew $(\sigma,a)$-constacyclic codes equals the number of different $(\mathbf{n},\sigma)$-isometry classes of skew $(\sigma,a)$-constacyclic codes.

### v3

@@ 5.3. An interesting correction
Let $\sigma\in{\rm Aut}(S)$ have order $\mathbf{m}$, fixed field $S_{0}$, and let $a,b,\alpha\in S^{\times}$. If there is a Hamming-weight preserving isomorphism $G:\mathbb{S}_{t^{\mathbf{n}}-a}\to\mathbb{S}_{t^{\mathbf{n}}-b}$ with $G|_{S}=\tau$, where $\tau\in\operatorname{Aut}(S)$ commutes with $\sigma$, then
@@
Suppose there exists a Hamming-weight preserving ring isomorphism given by $G:\mathbb{S}_{t^{\mathbf{n}}-a}\to\mathbb{S}_{t^{\mathbf{n}}-b}$ such that $G|_{S}=\tau$ and $G(t)=\alpha t^{k}$ for some $\alpha\in S^{\times}$, $0<k\leq\mathbf{n}-1$.
@@
    Suppose $G_{\tau,\alpha,k}:\mathbb{S}_{t^{\mathbf{n}}-a}\to\mathbb{S}_{t^{\mathbf{n}}-b}$ is an isomorphism with $k>1$ for some $\alpha\in S^{\times}$. Then $k\equiv 1\mod\mathbf{m}$, $\gcd(k,\mathbf{n})=1$, and $G_{\tau,\alpha}:\mathbb{S}_{t^{\mathbf{n}}-a}\to\mathbb{S}_{t^{\mathbf{n}}-b^{k}}$ is also an isomorphism.
@@
    Then $G_{\tau,\alpha}:\mathbb{S}_{t^{\mathbf{n}}-a}\to\mathbb{S}_{t^{\mathbf{n}}-b}$ is an isomorphism if and only if $G_{id,\alpha}:\mathbb{S}_{t^{\mathbf{n}}-\tau(a)}\to\mathbb{S}_{t^{\mathbf{n}}-b}$ is an isomorphism.
@@
    Also, for any $a\in\mathrm{Fix}(\tau)=\{b\in S:\tau(b)=b\}$, we have $G_{\operatorname{id},\alpha}:\mathbb{S}_{t^{\mathbf{n}}-a}\to\mathbb{S}_{t^{\mathbf{n}}-b}$ is an isomorphism if and only if $G_{\tau,\alpha}:\mathbb{S}_{t^{\mathbf{n}}-a}\to\mathbb{S}_{t^{\mathbf{n}}-b}$ is an isomorphism.
@@
$G_{\tau,\alpha,k}:\mathbb{S}_{t^{\mathbf{n}}-a}\to\mathbb{S}_{t^{\mathbf{n}}-b}$
@@
$G_{\tau,\alpha}:\mathbb{S}_{t^{\mathbf{n}}-a}\to\mathbb{S}_{t^{\mathbf{n}}-b^{k}}$
@@
$G_{id,\alpha}:\mathbb{S}_{t^{\mathbf{n}}-\tau(a)}\to\mathbb{S}_{t^{\mathbf{n}}-b}$
being a ring isomorphism that restricts to the identity on $S_{0}$.
@@
Moreover, (a) implies that two classes of skew constacyclic codes that are $(\mathbf{n},\sigma)$-isometric and related to principal left ideals in $\mathbb{S}_{t^{\mathbf{n}}-a}$, respectively to principal left ideals in $\mathbb{S}_{t^{\mathbf{n}}-b}$, will generate two classes of codes in $\mathbb{S}_{t^{\mathbf{n}}-a}$, respectively, in $\mathbb{S}_{t^{\mathbf{n}}-b}$, that are $(m,\sigma)$-equivalent.
@@
We derive the answer for the case that $\mathbb{S}_{t^{\mathbf{n}}-a}=S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-a)$ is nonassociative, that is, when $(\mathbf{n},\sigma)$-equivalence and $(\mathbf{n},\sigma)$-isometry coincide. Note that the proof does not require that $t^{n}-a$ is reducible.
@@
Suppose $K=\mathbb{F}_{p^{r}}$ and $\sigma(x)=x^{p^{s}}$ with $s|r$ so that $\mathbf{m}=r/s$ and $K_{0}=\mathbb{F}_{p^{s}}$. The number of distinct $(\mathbf{n},\sigma)$-isometry classes of families of skew $(\sigma,a)$-constacyclic codes arising from nonassociative rings $K[t;\sigma]/K[t;\sigma](t^{\mathbf{n}}-a)$ is $N$ where
$N=\begin{cases}\gcd([\mathbf{n}]_{s},p^{r}-1)&\text{if $\mathbf{m}\nmid\mathbf{n}$; and}\\ \left(1-\frac{1}{[\mathbf{m}]_{s}}\right)\gcd([\mathbf{n}]_{s},p^{r}-1)&\text{if $\mathbf{m}|\mathbf{n}$.}\end{cases}$
There are additionally $\gcd([\mathbf{n}]_{s},p^{r}-1)/[\mathbf{m}]_{s}$ different skew $(\mathbf{n},\sigma)$-equivalence classes (and thus at most this many $(\mathbf{n},\sigma)$-isometry classes) of families of skew $(\sigma,a)$-constacyclic codes arising from associative rings $K[t;\sigma]/K[t;\sigma](t^{\mathbf{n}}-a)$, that is, for which $\mathbf{m}\mid\mathbf{n}$ and $a\in K_{0}$.
@@
We start by recalling that the notions of $(\mathbf{n},\sigma)$-equivalence and $(\mathbf{n},\sigma)$-isometry coincide when $\mathbb{K}_{t^{\mathbf{n}}-b}:=K[t;\sigma]/K[t;\sigma](t^{\mathbf{n}}-b)$ is a proper nonassociative ring.
For any $a,b\in K$, the class of skew $(\sigma,a)$-constacyclic codes is $(\mathbf{n},\sigma)$-equivalent to the class of skew $(\sigma,b)$-constacyclic codes if and only if
@@
for some $\alpha\in K^{\times}$, that is, if and only if $a$ and $b$ lie in the same coset of the subgroup $N_{n}^{\sigma}(K^{\times})$. Let $\xi$ be a primitive element of $K$; it has order $p^{r}-1$. Then $N_{\mathbf{n}}^{\sigma}(K^{\times})$ is generated by $N_{\mathbf{n}}^{\sigma}(\xi)=\xi^{[\mathbf{n}]_{s}}=\xi^{w}$ with $w=\gcd([\mathbf{n}]_{s},p^{r}-1)$, and thus the order of this group is
@@
First suppose that $\mathbf{m}\nmid\mathbf{n}$, so that all of the rings $\mathbb{K}_{t^{\mathbf{n}}-a}$ are not associative. In this case, every coset $aN_{\mathbf{n}}^{\sigma}(K^{\times})$ corresponds to a class of $(\sigma,a)$-constacyclic codes such that $\mathbb{K}_{t^{\mathbf{n}}-a}$ is not associative, and there are
$|K|/|N_{\mathbf{n}}^{\sigma}(K^{\times})|=\gcd([\mathbf{n}]_{s},p^{r}-1)=w$
such classes.
Now suppose $\mathbf{m}|\mathbf{n}$, so that we are in the nonassociative case if and only if $a,b\in K\setminus K_{0}$. Recall $\mathbf{m}=r/s$. Then $N_{\mathbf{n}}^{\sigma}(\mathbb{F}_{p^{r}})\subset\mathbb{F}_{p^{s}}$, so that every coset of $N_{\mathbf{n}}^{\sigma}(\mathbb{F}_{p^{r}})$ is either entirely contained in $\mathbb{F}_{p^{s}}$ or entirely disjoint from it (so corresponds to a nonassociative Petit ring).
Note that $\mathbf{m}|\mathbf{n}$ implies $r|s\mathbf{n}$ so that $p^{r}-1$ divides $p^{s\mathbf{n}}-1$. Since $\mathbf{n}]_{s}=\frac{p^{s\mathbf{n}}-1}{p^{s}-1}$ we deduce that $\mathbf{m}]_{s}=(p^{r}-1)/(p^{s}-1)$ divides $w=\gcd([\mathbf{n}]_{s},p^{r}-1)$. With this in mind we now separate the two cases.
The total number of cosets contained in $\mathbb{F}_{p^{s}}$, and thus corresponding to associative rings $K[t;\sigma]/K[t;\sigma](t^{\mathbf{n}}-a)$ with $a\in\mathbb{F}_{p^{s}}$, is equal to
$|\mathbb{F}_{p^{s}}^{\times}|/|N_{\mathbf{n}}^{\sigma}(K^{\times})|=\frac{p^{s}-1}{(p^{r}-1)/w}=\frac{w}{[\mathbf{m}]_{s}}\in\mathbb{N}.$
In that case, we have shown that the notions of $(\mathbf{n},\sigma)$-isometry and $(\mathbf{n},\sigma)$-equivalence need not coincide. Thus these $(\mathbf{n},\sigma)$-isometry classes may be bigger (thus fewer in number), yielding the last sentence of the theorem.
Finally, subtracting these $\frac{w}{[\mathbf{m}]_{s}}$ cosets from the total number of cosets gives the number of $(\mathbf{n},\sigma)$-isometry classes for nonassociative rings in the case that $\mathbf{m}|\mathbf{n}$, that is, that $N=w-w/[\mathbf{m}]_{s}$, as required. ∎
This recovers, in particular, [OaNA25, Theorem 6] as a special case, as this means that the number of $(\mathbf{n},\sigma)$-equivalence classes is $\gcd([\mathbf{n}]_{s},p^{r}-1)$.
@@
At one extreme, suppose that $N_{\mathbf{n}}^{\sigma}(S^{\times})=S^{\times}$. Then all of the rings $\mathbb{S}_{t^{\mathbf{n}}-a}=S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-a)$ are $(\mathbf{n},\sigma)$-equivalent, and in particular all are $(\mathbf{n},\sigma)$-equivalent to $\mathbb{S}_{t^{\mathbf{n}}-1}$ which corresponds to the class of skew cyclic codes. Moreover, equivalence and $(\mathbf{n},\sigma)$-equivalence coincide in this case. When $\mathbf{m}\nmid\mathbf{n}$, then the notions of equivalence, isometry and $(\mathbf{n},\sigma)$-isometry coincide as well. Hence for all $a\in S^{\times}$, the skew $(\sigma,a)$-constacyclic codes of length $\mathbf{n}$ over $S$ are equivalent to the skew cyclic codes of length $\mathbf{n}$.
As a special case, let $K=\mathbb{F}_{p^{r}}$, $\sigma(x)=x^{p^{s}}$ with $s|r$ and $K_{0}=\mathbb{F}_{p^{s}}$. If $\gcd([\mathbf{n}]_{s},p^{r}-1)=1$ then $N_{\mathbf{n}}^{\sigma}(K^{\times})=K^{\times}$ so all $K[t;\sigma]/K[t;\sigma](t^{\mathbf{n}}-a)$ are $(\mathbf{n},\sigma)$-equivalent, as was already observed in [OaNA25, Corollary 2] and [BBB21, Proposition 1].
@@
At the other extreme, when $\gcd([\mathbf{n}]_{s},p^{r}-1)=p^{r}-1$ (as in Example 5.5), then $N_{\mathbf{n}}^{\sigma}(K^{\times})=\{1\}$ and so no two distinct $a,b\in K^{\times}$ will be $(\mathbf{n},\sigma)$-equivalent. Thus, there are a plethora of distinct classes of skew constacyclic codes up to $(\mathbf{n},\sigma)$-equivalence. That said, there always $r$ choices for $\tau\in{\rm Gal}(\mathbb{F}_{p^{r}}/\mathbb{F}_{p})$.
@@
On the other hand, when $\mathbb{S}_{t^{\mathbf{n}}-a}$ is associative we can prove a powerful equivalence (inspired by the flavour of [OaNA25, Theorem 4]) that holds for any commutative ring $S$ and any $\tau$ commuting with $\sigma$.
@@
    $S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-a)$ and $S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$ are isometric via $G_{\tau,\alpha,k}$;
@@
    $S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b^{k})$ and $S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-a)$ are equivalent via $G_{\tau,\alpha}$, where $k$ satisfies $k\equiv 1\mod\mathbf{m}$ and $\gcd(k,\mathbf{n})=1$.